OpenStudy (anonymous):

2x + 3y = 2 x + 6y = 4 help

OpenStudy (anonymous):

What are you trying to do?

OpenStudy (anonymous):

solve for (x) and (y)

OpenStudy (anonymous):

First, isolate one of the variables, in this case, x. 2x + 3y = 2 ----> 2x = -3y + 2 ---> x = -3y/2 + 1 Then, plug that into the second equation and solve for the other variable. (-3y/2 + 1) + 6y = 4 Solve for y. -3y + 6y = 3/2 Combine like terms. 3y=3/2 y=1/2 Finally, plug that y value into either equation and solve for x.

OpenStudy (anonymous):

so it would be x=4 and y=2

OpenStudy (anonymous):

no x=4/3 right @gatorsfan75

OpenStudy (anonymous):

It would be y= 1/2 and x=1, are those answer choices?

OpenStudy (anonymous):

no my choices are x=0 , y= 2/3 x=2 , y= -2/3 x=4/3 , y = 2/9 x=4 , y= -2

OpenStudy (anonymous):

@gatorsfan75

OpenStudy (anonymous):

Then the answer is x=0 and y=2/3

OpenStudy (anonymous):

Plug them into both equations to double check my work :)

OpenStudy (anonymous):

2x + 3y =2 -------------> (2x0=0) + (3x2/3= um im not sure how to do that @gatorsfan75

OpenStudy (anonymous):

2(0) + 3(2/3) = 2 0 + 2 = 2 (CORRECT) Next equation: (0) + 6(2/3) = 4 0 + 4 = 4 (CORRECT)

OpenStudy (anonymous):

Those are your answers :)

OpenStudy (anonymous):

thanks alot @gatorsfan75